# How Haag-tied is QFT, really?

### 18 Dec 2023 v2

@@ 4.2 Theory Interpreters—The Particle Problem
Thus, for these authors, Haag’s theorem sits alongside results like the Reeh–Schlieder theorem [Reeh and Schlieder, 1961], the Unruh effect ([Fulling, 1973], [Davies, 1975], [Unruh, 1976] and discussed more recently in [Crispino et al., 2008] [Earman, 2011]), or the Hegerfeldt [Hegerfeldt, 1998a] [Hegerfeldt, 1998b] and Malament [Malament, 1996] theorems (and extensions thereof [Halvorson and Clifton, 2002]) as evidence against the existence of a fundamental particle interpretation for RQFT. Indeed, canon seems to dictate that these results be discussed together when addressing the particle problem (e.g., [Bain, 2000, 380] [Halvorson and Clifton, 2002, 20] [Fraser, 2008, 842] [Ruetsche, 2011, Chs. 9–11] [Wallace, 2001, §2.4].)
Crucially, the existence of a global number operator is necessary for the standard fundamental particle interpretation since these operators enforce the presumed right notion of localizability and countability inherent in the particle concept. For pristine interpreters, there is a further desideratum: the selfsame particle concept must be available throughout each timeslice of a model of particle interactions. Put so crudely, it is this latter aspect of the criterion that seems to rule out emergence- and coalescence-type particle notions like that found in [Wallace, 2001]. [^25] See also [Feintzeig et al., 2021] for recent work on accounting for emergent particle phenomena using the classical limit of QFT.
@@
Interpreters addressing the highly constrained puzzle of developing a particle interpretation of interacting QFT bite the bullet and accept that there are no fundamental particles. Doing so entails providing an alternative fundamental ontology and, consequently, alternative explanations of erstwhile “particle” physics phenomena. Halvorson and Clifton conclude that “relativistic quantum field theory does permit talk about particles—albeit, if we understand this talk as really being about the properties of, and interactions among, quantized fields” [Halvorson and Clifton, 2002, 24]. A large part of the repair offered in [Halvorson and Clifton, 2002] is an account of how our talk about ‘particles’ is still useful, despite its fictitious status. Similarly, albeit with less patience for the talk of particles, D. Fraser concludes that special relativity and the non-linearity of the field equation for an interacting system—key ingredients for Haag’s theorem—conspire against a quanta interpretation and lending some weak support to a field interpretation instead (“At least on the surface QFT is a theory of fields” and “there is no quanta interpretation and there are no quanta” [Fraser, 2008, 857-8].  [^26] However, it should be noted that the most obvious way of constructing a field ontology appears to run into the same problems as the particle ontology [Baker, 2009]. Baker’s argument relies on the restriction that the field wave functionals are square-integrable. If we relax this restriction, then the space of wave functionals will be larger than the space of particle wave functions (see [Jackiw, 1990] and [Sebens, 2022], but see [Wallace, 2006] for a defense of the square-integrability restriction). We may think of the repair work advocated here as follows. It is as though the ‘owner’ of our QFT building had intended to furnish the interacting QFT ‘room’ with a fundamental particle ontology to account for particle-like phenomena; this group of subcontracts is saying that the deep structure of the building simply will not allow for that kind of ‘furniture’ in that particular ‘room.’ The best we can do is paint some faux particles on the walls to save the usefulness of particle-talk.
If, however, we set aside the highly constrained puzzle within the program of rigorous formulations of QFT, we can pursue alternative programs of particle interpretations rooted in revised formal notions of ‘particle.’ These alternative programs can be pursued according to either pristine or non-pristine interpretive standards, but in either case one must revise the relevant definitions at play in the no-go on particles. One such option revises the basic approach to scattering theory (Haag–Ruelle) [Reed and Simon, 1975]. Another option revises the condition for localizability and state equivalence (LSZ formalism) [Bain, 2000]. Or we might instead consider an emergent notion of particles in QFT. Instead of proposing repairs to either our ontology or QFT, it suffices for [Wallace, 2001] to show that a concept of particle can be recovered from a field-theoretic description. Given Wallace’s background presumption that approximate concepts are legitimate in physics, it does not matter that the particle concept he deploys is vaguely defined and applies only to bosonic, massive QFTs; likewise, the satisfaction with approximate concepts removes the sting of Haag’s theorem for a fundamental particle notion. For Ruetsche, Haag’s theorem is only one chapter (namely, Ch. 9) in the story of particles [Ruetsche, 2011, Chs. 9–11], and that it rules out the standard particle interpretation in general is unremarkable in the larger story. Indeed, she concludes, “[s]ometimes particle physics is, adulteratedly, particle physics, and that’s a good thing” [Ruetsche, 2011, 260].
@@
Non-pristine interpreters, we can safely assume, set their long term sights on increasing the thoroughness of their multiple, context-dependent interpretations of QFT. Haag’s theorem itself is not an especially valuable guide to such a project. The long-term desire of the pristine interpreter group—including each of the above repairs—is to further develop and refine a pristine interpretation of QFT. Given this group’s expectations for how one should be developed, a mathematically precise, general, and consistent formulation of QFT, both in its foundations and in specific models, is a goal. And under the auspices of pristine interpretation, Haag’s theorem is a key guidepost. Whereas Duncan and Miller satisfy themselves with identifying violations of the assumptions of Haag’s theorem in particular settings, thus potentially making available in those settings a particle interpretation amenable to the non-pristine interpreter, the pristine interpreter requires a satisfactory response applicable across all settings. This feeds directly into the resource allocations proposed by pristine interpreters: since a response must work across all settings, axiomatic approaches to inquiry are naturally preferred. For this reason, satisfying responses are typically sought from the formal variants of QFT; the effective field theory approach, in particular, is not the right place to look for long term stability. Thus, expected renovations for this group are largely mathematical. We note here that current mathematically rigorous models of QFT are deficient in various ways, such as the decades long failed search for an interacting model in four spacetime dimensions. [^27] Recall D. Fraser’s assessment of this search in note 14. One noteworthy recent development on what kinds of field theories can be rigorously constructed in four space-time dimensions is [Aizenman and Duminil-Copin, 2021]. Likewise, revisions of the relevant definitions (e.g., Bain’s re-definition of ‘particle’) have led to difficulties [Fraser, 2008]. Long term successful renovations in this area will likely require deep conceptual work in tandem with mathematical precise explications of those concepts.
Another line of renovation work is advocated in [Bain, 2011]. He argues that our intuitions about what formal requirements capture the appropriate senses of localizability and countability constitutive of the particle notion at stake in these debates are themselves non-relativistic. Holding a putative particle interpretation of relativistic QFT to non-relativistic standards of localizability and countability is misguided. Instead, Bain urges, renovation efforts ought to center on developing an appropriately relativistic representation of localizability and countability. Similarly, projects aimed at explicating an emergent notion of particles ([Wallace, 2001], [Feintzeig et al., 2021]) advocate long term renovation work in terms of conceptual criteria for particle interpretations.
@@ 4.3 Duncan and Miller
Tony Duncan [Duncan, 2012] and Michael Miller [Miller, 2018] argue that at least one of the assumptions needed to prove Haag’s theorem is violated at some point in the actual process of calculating scattering theory results in perturbative QFT. Thus, in Duncan’s words, we can “stop worrying” about Haag’s theorem. These calculational processes include the methods of regularization and renormalization.  [^28] Regularization is a mathematical procedure used to temporarily control the (UV or IR) divergences that arise in QFT calculations. After regularization, it becomes possible to extract finite results from the perturbative series, order-by-order. However any regularization technique will introduce new, arbitrary parameters (“regularization scales”). Renormalization is a subsequent procedure in which the arbitrary dependence on these regularization scales is absorbed by redefining the bare parameters of the theory. For a fuller explanation, see [Collins, 1984].
A similar view on Haag’s theorem is given in [Fraser, 2017]. Here, James Fraser is primarily concerned with the much broader question of the status of perturbative QFT and diagnosing the ‘real problem’ with this area of physics. He argues that perturbative QFT is “a method for producing approximations without addressing the project of constructing interacting QFT models” (4).  [^29] Miller also addresses the question of what happens after we remove UV and IR cutoffs and thus restore exact Poincare covariance, which J. Fraser does not directly discuss. As a consequence of this view, the threat of inconsistency posed by Haag’s theorem is defused for much the same reasons as given by Duncan and Miller. J. Fraser concludes section 4, “The perturbative method simply does not assert the set of claims shown to be inconsistent by Haag’s theorem” (18). We set aside [Fraser, 2017] for the remainder of this section since Haag’s theorem is not the primary target of that article; but readers interested in the broader question of how to assess perturbative QFT are encouraged to look there for an important contribution on that topic.
@@
Both Duncan and Miller recognize the logical nature of the problem posed by Haag’s theorem—i.e., as the conflict between the interaction picture and free-field representation assumptions—and they take consistency to be a requirement for a viable theory. Indeed, they recognize that a response to Haag’s theorem must come from negating at least one of its assumptions. However, unlike Earman and D. Fraser, their goal is to “explain why theoretical predictions for realistic experimental observables give empirically adequate results” [Miller, 2018, 803]. These predictions in fact apply (a version of) the interaction picture, so they cannot simply jettison the framework of the interaction picture altogether**. Rather, they aim to identify where within the practical calculational techniques such violations of the assumptions must already take place for reasons entirely independent of Haag’s theorem.
@@
The interaction picture in practice makes use of renormalization and regularization techniques that, in pulling “double duty”, both produce finite perturbative results and defuse the assumptions of Haag’s theorem. The renormalization and regularization ‘repairs’ also evade Haag’s theorem as a positive side effect. No additional repair is needed.
@@
Maintenance is straightforward: maintain current practices of proper use of renormalization and regularization techniques. For the conceptually and logically curious, following in Duncan and Miller’s footsteps, this maintenance will likely also involve identifying precisely which assumptions have been violated, and where. While it remains to be shown that the perturbative power-series converge and thus correspond to exact non-perturbative objects post-regularization and -renormalization, [Miller, 2018, page 815] contends that this problem is both unrelated to Haag’s theorem and more serious (compare [Miller, 2021, 1135]).
@@ 4.4 Klaczynski
Given the above, Klaczynski approaches Haag’s theorem intent on understanding what it says about canonical QFT. At first blush, the result appears to be negative. For instance, the Gell-Mann-Low formula, relating the ground states of the interacting and non-interacting fields, is built on the assumption that the time-evolution operator in the interaction picture, which relates the two fields, is unitary; this is exactly what Haag’s theorem rules out (recall note 7 on the significance and wide-spread use of the Gell-Mann-Low formula.). However, on closer inspection, the contradiction is resolved, if nevertheless unfavorably: like J. Fraser and Duncan, Klaczynski blames Haag’s theorem for the divergence of the perturbative expansion of the Gell-Mann-Low formula. This leads Klaczynski to conclude that the interaction picture is ill-defined and trivial [Klaczynski, 2016, 59].
@@
Like Duncan and Miller, Klaczynski thinks that renormalization procedure still repairs the problem, but gives a different reason for why it works. When we renormalize our theories, we replace the bare quantities with their renormalised counterparts. This process of renormalization does not merely cancel the divergences, but also fundamentally changes the theory, by bringing about a coupling-dependent mass shift. As a result, the renormalized free and interacting fields now have different masses. Two quantum fields of different mass are overwhelmingly likely to be unitarily inequivalent at any given time. 323232 The mathematical ill-definedness of the renormalized terms makes it hard to prove that renormalized free and interacting fields will be unitarily inequivalent in general. However, Klaczynski uses a theorem X.46 of [Reed and Simon, 1975] to argue that it is “plausible beyond doubt” [Klaczynski, 2016, page 68]. Theorem X.46 says that if there are two free, scalar, neutral fields with a unitary intertwiner at some time $t_{0}$, with different masses, then these fields must be unitarily inequivalent. Klaczynski uses the theorem to argue we should not expect a unitary intertwiner at any time between the renormalized interacting and free fields. Recall that Haag’s theorem requires precisely such a unitary equivalence between the free and interacting fields at some time (10). , and so the conditions for Haag’s theorem do not apply in our renormalized theories.
@@
Given that the interaction-picture presentation of canonical QFT still dominates, renovation is necessary. This has two parts. First, physicists have misunderstood, or at least mis-described, the tools they are using when performing scattering calculations with renormalized fields. While Haag’s theorem says this tool cannot be the interaction picture, that does not mean renormalization is unconstrained or incoherent. Rather, renormalization “follows rules which have a neat underlying algebraic structure [the Hopf algebra] and are not those of a random whack-a-mole game” [Klaczynski, 2016, page 4].
Second, this program of identifying the structure of renormalization must continue. First and foremost, there are some mathematical lacunae in this process, as well as procedures that are not defined wholly rigorously; these should be addressed. Klaczynski believes that renormalized quantum field theory “provides us with peepholes through which we are allowed to glimpse at least some parts of that ‘true’ theory” [Klaczynski, 2016, page 4]. As such, the ultimate goal here would be to glimpse what we can of the ‘true’ theory through the peepholes this work affords.  [^33] This view can be fruitfully compared to that of [Fraser, 2017]. Both agree that the perturbative renormalization offers a consistent calculational procedure that can both generate empirical predictions and offers some insights about underlying “true” physics. However, J. Fraser argues for a selectively realist attitude towards some parts of the effective theory. Klaczynski seems to argue only that this theory might provide some insights about a “true” physical theory, one which gives a true representation of the physical system.

### 06 Apr 2024 v3

@@ 4.2 Theory Interpreters—The Particle Problem
Thus, for these authors, Haag’s theorem sits alongside results like the Reeh-Schlieder theorem [Reeh and Schlieder, 1961], the Unruh effect ([Fulling, 1973], [Davies, 1975], [Unruh, 1976] and discussed more recently in [Crispino et al., 2008] [Earman, 2011]), or the Hegerfeldt [Hegerfeldt, 1998a] [Hegerfeldt, 1998b] and Malament [Malament, 1996] theorems (and extensions thereof [Halvorson and Clifton, 2002]) as evidence against a particle interpretation for RQFT. Indeed, canon seems to dictate that these results be discussed together when addressing the particle problem (e.g., [Bain, 2000, 380] [Halvorson and Clifton, 2002, 20] [Fraser, 2008, 842] [Ruetsche, 2011, Chs. 9–11] [Wallace, 2001, §2.4].)
@@
Interpreters addressing the highly-constrained puzzle of constructing a number operator for interacting QFT bite the bullet and accept that there are no fundamental particles. Doing so entails providing an alternative fundamental ontology and alternative explanations of erstwhile “particle” physics phenomena. Halvorson and Clifton conclude that “relativistic quantum field theory does permit talk about particles—albeit, if we understand this talk as really being about the properties of, and interactions among, quantized fields” [Halvorson and Clifton, 2002, 24]. A large part of the repair offered in [Halvorson and Clifton, 2002] is an account of how our talk about “particles” is still useful, despite its fictitious status. Similarly, albeit with less patience for the talk of particles, D. Fraser concludes that special relativity and the non-linearity of the field equation for an interacting system—key ingredients for Haag’s theorem—conspire against a quanta interpretation, lending some weak support to a field interpretation instead (“At least on the surface QFT is a theory of fields” and “there is no quanta interpretation and there are no quanta” [Fraser, 2008, 857-8]. [^24] However, it should be noted that the most obvious way of constructing a field ontology appears to run into the same problems as the particle ontology [Baker, 2009]. Baker’s argument relies on the restriction that the field wave functionals are square-integrable. If we relax this restriction, then the space of wave functionals will be larger than the space of particle wave functions (see [Jackiw, 1990] and [Sebens, 2022], but see [Wallace, 2006] for a defense of the square-integrability restriction).
If, however, we set aside the highly-constrained puzzle within the program of rigorous formulations of QFT, we can pursue alternative programs of particle interpretations rooted in revised formal notions of ‘particle.’ These alternative programs can be pursued according to various standards of theory interpretation, but in any case one must revise the relevant definitions at play in the no-go arguments for particles. One such option revises the basic approach to scattering theory (Haag-Ruelle) [Reed and Simon, 1975]. In this approach, one jettisons the interaction picture, and instead finds surrogates for interacting states among free states. Thus, the challenges of accessing a Fock space representation for the interacting theory disappear. Another option, the LSZ formalism, revises the relations between the interacting field and the asymptotic free fields. As shown in [Bain, 2000], the LSZ formalism defuses the argument that the interacting field must be free by replacing strong convergence with weak convergence. Thus, Bain argues that the LSZ formalism solves the conceptual problem posed by Haag’s theorem. Bain formulates asymptotic raising and lowering operators by taking appropriate limits of the interacting raising and lowering operators. These asymptotic operators are then used to define asymptotic particle states. The asymptotic particles states are not free field states, and so in principle there are contributions from the multi-particle continuum. Bain (§4.2) shows that these contributions are negligible on the relevant time scales, and thus, for all practical purposes, the asymptotic particle states replicate the behavior of free particle states. On the LSZ approach, then, Haag’s theorem is circumvented through the move to weak convergence; and the expected phenomenological difference between interacting states and free-like asymptotic states is recovered. The result is a rehabilitation of a particle interpretation through the redefinition of asymptotic states.
Or we might instead consider an emergent notion of particles in QFT. Instead of proposing repairs to either our ontology or the assumptions of Haag’s theorem, for [Wallace, 2001] it suffices to show that a concept of particle can be recovered from a field-theoretic description. Given Wallace’s background presumption that approximate concepts are legitimate in physics, it does not matter (as much) that the particle concept he deploys is vaguely defined and applies only to bosonic, massive QFTs.
[…]
@@
Here theory interpreters especially diverge from the rest of our authors. For both those attendant to the highly-constrained puzzle —given QFT in a mathematically rigorous form, is it possible to construct a number operator for interacting QFT?—and those who are not, the interpretive projects seek to establish an ontology conditional on an extant form of QFT. This interpretive task is a step removed from questions about how to develop or innovate with our theory. As such, the renovations that theory interpreters propose to improve interpretation will be mediated by how they think the theory should be developed. Given that this latter question falls outside the domain of theory interpretation per se, theory interpreters often say comparatively little about what renovations are necessary in response to Haag’s theorem.
Nevertheless, we can say a little bit about what renovations theory interpreters might call for. For those trying to solve the puzzle, we need a rigorous formulation of QFT or its successor which we can subject to interpretation. To answer the puzzle, this formulation must be general enough that we can draw interpretive conclusions about all of the phenomena we consider relevant, not just a subset. These calls would seem to make puzzle solvers inclined to axiomatic approaches to theory renovation and, in general, to development of formal variants of QFT. [^25] Recall D. Fraser’s assessment of the search for a mathematically rigorous model of QFT in note 14. One noteworthy recent development on what kinds of field theories can be rigorously constructed in four spacetime dimensions is [Aizenman and Duminil-Copin, 2021]. On the other hand, those less dedicated to the original puzzle seem to call for renovations of the basic conceptual apparatus, especially the notion of particle. [Bain, 2011] argues that the intuitions of localizability and countability constitutive of the particle notion at stake in these debates are themselves non-relativistic, and urges us to develop one more appropriate for relativistic theories. Similarly, interpretive projects aimed at explicating an emergent notion of particles ([Wallace, 2001], [Feintzeig et al., 2021]) advocate long term renovation work in terms of conceptual criteria for particle interpretations.
@@ 4.3 Duncan and Miller
Tony Duncan [Duncan, 2012] and Michael Miller [Miller, 2018] argue that at least one of the assumptions needed to prove Haag’s theorem is violated at some point in the actual process of calculating scattering theory results in perturbative QFT. Thus, in Duncan’s words, we can “stop worrying” about Haag’s theorem. These calculational processes include the methods of regularization and renormalization.  [^26] Regularization is a mathematical procedure used to temporarily control the ultraviolet (UV) or infrared (IR) divergences that arise in QFT calculations. After regularization, it becomes possible to extract finite results from the perturbative series, order-by-order. However any regularization technique will introduce new, arbitrary parameters (regularization scales). Renormalization is a subsequent procedure in which the arbitrary dependence on these regularization scales is absorbed by redefining the bare parameters of the theory. For a fuller explanation, see [Collins, 1984].
A similar view on Haag’s theorem is given in [Fraser, 2017]. Here, James Fraser is primarily concerned with the much broader question of the status of perturbative QFT and diagnosing the “real problem” with this area of physics. He argues that perturbative QFT is “a method for producing approximations without addressing the project of constructing interacting QFT models” (4).  [^27] Miller also addresses the question of what happens after we remove UV and IR cutoffs and thus restore exact Poincare covariance, which J. Fraser does not directly discuss. As a consequence of this view, the threat of inconsistency posed by Haag’s theorem is defused for much the same reasons as given by Duncan and Miller. J. Fraser concludes section 4, “The perturbative method simply does not assert the set of claims shown to be inconsistent by Haag’s theorem” (18). We set aside [Fraser, 2017] for the remainder of this section since Haag’s theorem is not the primary target of that article; but readers interested in the broader question of how to assess perturbative QFT are encouraged to look there for an important contribution on that topic.
@@
Both Duncan and Miller recognize the logical nature of the problem posed by Haag’s theorem—i.e., as the conflict between the interaction picture and free-field representation assumptions—and they take consistency to be a requirement for a viable theory. Indeed, they recognize that a response to Haag’s theorem must come from negating at least one of its assumptions. However, unlike Earman and D. Fraser, their goal is to “explain why theoretical predictions for realistic experimental observables give empirically adequate results” [Miller, 2018, 803]. These predictions in fact apply (a version of) the interaction picture, so they cannot simply jettison the framework of the interaction picture altogether. Rather, they aim to identify where within the practical calculational techniques such violations of the assumptions must already take place for some other reasons unmotivated by Haag’s theorem.
@@
The interaction picture in practice makes use of renormalization and regularization techniques that, in pulling “double duty”, both produce finite perturbative results and defuse the assumptions of Haag’s theorem. The renormalization and regularization repairs also evade Haag’s theorem as a positive side effect. No additional repair is needed.
@@
Maintenance is straightforward: maintain current practices of proper use of renormalization and regularization techniques. For the conceptually and logically curious, following in Duncan and Miller’s footsteps, this maintenance will likely also involve identifying precisely which assumptions have been violated, and where. While it remains to be shown that the perturbative power series converge and thus correspond to exact non-perturbative objects post-regularization and -renormalization, [Miller, 2018, page 815] contends that this problem, while more serious, is ultimately unrelated to Haag’s theorem (compare [Miller, 2021, 1135]).
@@ 4.4 Klaczynski
Given the above, Klaczynski approaches Haag’s theorem intent on understanding what it says about canonical QFT. At first blush, the result appears to be negative. For instance, the Gell-Mann and Low formula, relating the ground states of the interacting and non-interacting fields, is built on the assumption that the time-evolution operator in the interaction picture, which relates the two fields, is unitary; this is exactly what Haag’s theorem rules out (recall note 7 on the significance and widespread use of the Gell-Mann and Low formula.). However, on closer inspection, the contradiction is resolved, if nevertheless unfavorably: like J. Fraser and Duncan, Klaczynski blames Haag’s theorem for the divergence of the perturbative expansion of the Gell-Mann and Low formula. This leads Klaczynski to conclude that the interaction picture is ill-defined and trivial [Klaczynski, 2016, 59].
@@
Like Duncan and Miller, Klaczynski thinks that renormalization procedure still repairs the problem, but gives a different reason for why it works. When we renormalize our theories, we replace the bare quantities with their renormalised counterparts. This process of renormalization does not merely cancel the divergences, but also fundamentally changes the theory, by bringing about a coupling-dependent mass shift. As a result, the renormalized free and interacting fields now have different masses. Two quantum fields of different mass are overwhelmingly likely to be unitarily inequivalent at any given time. 303030 The mathematical ill-definedness of the renormalized terms makes it hard to prove that renormalized free and interacting fields will be unitarily inequivalent in general. However, Klaczynski uses a theorem X.46 of [Reed and Simon, 1975] to argue that it is “plausible beyond doubt” [Klaczynski, 2016, page 68]. Theorem X.46 says that if there are two free, scalar, neutral fields with a unitary intertwiner at some time $t_{0}$, with different masses, then these fields must be unitarily inequivalent. Klaczynski uses the theorem to argue we should not expect a unitary intertwiner at any time between the renormalized interacting and free fields. Recall that Haag’s theorem requires precisely such a unitary equivalence between the free and interacting fields at some time (10). , and so the conditions for Haag’s theorem do not apply in our renormalized theories.
@@
Given that the interaction picture presentation of canonical QFT still dominates, renovation is necessary. This has two parts. First, physicists have misunderstood, or at least misdescribed, the tools they are using when performing scattering calculations with renormalized fields. While Haag’s theorem says this tool cannot be the interaction picture, that does not mean renormalization is unconstrained or incoherent. Rather, renormalization “follows rules which have a neat underlying algebraic structure [the Hopf algebra] and are not those of a random whack-a-mole game” [Klaczynski, 2016, page 4].
Second, this program of identifying the structure of renormalization must continue. First and foremost, there are some mathematical lacunae in this process, as well as procedures that are not defined wholly rigorously; these should be addressed. Klaczynski believes that renormalized quantum field theory “provides us with peepholes through which we are allowed to glimpse at least some parts of that ‘true’ theory’ ” [Klaczynski, 2016, page 4].  [^31] This view can be fruitfully compared to that of [Fraser, 2017]. Both agree that the perturbative renormalization offers a consistent calculational procedure that can both generate empirical predictions and offers some insights about underlying “true” physics. However, J. Fraser argues for a selectively realist attitude towards some parts of the effective theory. Klaczynski seems to argue only that this theory might provide some insights about a “true” physical theory, one which gives a true representation of the physical system.